OpenStudy (jessicajohnson):

sin 2 theta - cos theta =0

OpenStudy (anonymous):

what is theta?

OpenStudy (anonymous):

\[\sin 2\theta-\cos \theta=0\]

OpenStudy (jessicajohnson):

Thank you Raphael!!

OpenStudy (anonymous):

you are welcome:)